# Transmission Conditions for the Non-Overlapping Schwarz Coupling of Full Order and Operator Inference Models

### v1

@@ Abstract
This work investigates transmission conditions for the domain decomposition-based coupling of subdomain-local models using the non-overlapping Schwarz alternating method (NO-SAM). Building on prior efforts involving overlapping SAM (O-SAM), we formulate and assess two NO-SAM variants, based on alternating Dirichlet–Neumann and Robin–Robin transmission conditions, respectively. For the subdomain-local models, we consider a mix of full order models (FOMs) and non-intrusive reduced order models (ROMs) constructed via an emerging model reduction technique known as operator inference (OpInf). Of particular novelty is the first application of NO-SAM to couple non-intrusive OpInf ROMs with each other and with FOMs. Numerical studies on a one-dimensional linear elastic wave propagation benchmark problem demonstrate that transmission condition choice and parameter tuning significantly impact convergence rate, accuracy, and stability. Robin–Robin coupling often yields faster convergence than alternating Dirichlet–Neumann, though improper parameter selection can induce spurious oscillations at subdomain interfaces. For FOM–OpInf and OpInf–OpInf couplings, higher modal content in the ROM basis improves accuracy and mitigates instabilities, in some cases outperforming coupled FOM–FOM reference solutions when it comes to accuracy. These findings highlight NO-SAM’s potential for enabling flexible, non-intrusive, and efficient multi-model coupling across independently meshed subdomains, while emphasizing the need for careful interface condition design in higher-dimensional and predictive settings.
@@ 1 Introduction
The present work advances a computational technique that promises to mitigate both hurdles described above: the Schwarz alternating method (SAM) [39]. The key idea behind SAM is to decompose the physical domain on which a given partial differential equation (PDE) is posed into smaller subdomains, and then solve the governing PDE by iterating between smaller, subdomain-local problems, while exchanging boundary condition information to ensure compatibility across subdomain interfaces. Over the past decade, our group has demonstrated SAM’s ability to achieve concurrent domain decomposition- (DD-)based coupling of subdomain-local full order models (FOMs) with each other [33, 34, 32], as well as with data-driven ROMs of various flavors (e.g., intrusive projection-based ROMs [46, 1], physics-informed neural networks (PINNs) [42], non-intrusive operator inference (OpInf) models [31]). In these papers, we have demonstrated that SAM has a number of advantages over competing multi-scale coupling methods. Most notably: it is concurrent and minimally-intrusive, can couple non-conformal meshes with different element topologies, and allows each subdomain to be advanced with different time integrators and time steps for dynamic problems. Importantly, SAM achieves this without introducing non-physical artifacts into the solution.
@@
The work described here is an addition to a growing body of literature on DD-based OpInf-OpInf and ROM-FOM couplings. Our SAM approach is most similar to the recent work by Farcas et al. [13], which develops a method for performing DD-based coupling of subdomain-local OpInf ROMs by learning appropriate reduced operators responsible for the coupling. The present work differs from that of Farcas et al. in two important ways: (i) in [13], each subdomain problem is solved once rather than by performing an iteration to convergence as done within our Schwarz framework, which may be insufficient for complex nonlinear problems, and (ii) the approach in [13] is limited to overlapping domain decompositions (DDs). While other Schwarz-like methods for FOM-FOM and FOM-OpInf couplings exist in the literature, the majority of these methods focus on the coupling of intrusive projection-based ROMs [6, 24, 25, 38] or neural network- (NN-)based ROMs [44, 45, 27, 28, 42]. A variety of other FOM-FOM and FOM-OpInf coupling methods, e.g., Lagrange multipliers, optimization-based coupling, coupling via boundary maps, each with their own advantages and disadvantages, have also been proposed over the years. The interested reader is referred to [5, 7, 8, 23, 30, 36, 37, 11, 18, 3, 16, 41, 11, 19, 4, 43, 21, 12] and the references therein for more details. To the best of the authors’ knowledge, only two of these references, namely [16] and [13], consider the DD-based coupling of subdomain-local non-intrusive OpInf ROMs of the flavor considered herein [35].
@@ 2 Solid mechanics problem formulation
$\mathrm{Div}\,~P+\rho_{0}~B=\rho_{0}\ddot{\boldsymbol{\Psi}}\quad\text{in }\Omega\times I$ (1)
where $\Omega\in\mathbb{R}^{d}$ for $d\in\{1,2,3\}$ is an open bounded domain, $I\coloneqq\{t\in[t_{0},t_{f}]\}$ is a closed time interval with $0\leq t_{0}<t_{f}$, and $x=\boldsymbol{\Psi}(~X,t):\Omega\times I\rightarrow\mathbb{R}^{d}$ describes the mapping from the reference configuration to the current configuration with $X\in\Omega$. The overdot notation in $\ddot{\boldsymbol{\Psi}}$ denotes a time derivative such that $\dot{\boldsymbol{\Psi}}\coloneqq\frac{\partial\boldsymbol{\Psi}}{\partial t}$ and $\ddot{\boldsymbol{\Psi}}\coloneqq\frac{\partial^{2}\boldsymbol{\Psi}}{\partial t^{2}}$. The symbol P~P denotes the first Piola-Kirchoff stress, B~B is the specific body force, and $\rho_{0}$ is the mass density defined in the reference configuration. It is important to note that the constitutive model used to describe P~P is general and can include nonlinear models.
@@
$\boldsymbol{\Psi}(~X,t_{0})$ $=~X_{0}$ $\text{in }\Omega$ (2)
$\dot{\boldsymbol{\Psi}}(~X,t_{0})$ $=~v_{0}$ $\text{in }\Omega$
$\boldsymbol{\Psi}(~X,t)$ $=\boldsymbol{\chi}$ $\text{on }\partial_{\Psi}\Omega\times I$
$~P~N$ $=~T$ $\text{on }\partial_{T}\Omega\times I$
where $\partial_{\Psi}\Omega$ and $\partial_{T}\Omega$ denote the Dirichlet and Neumann portions, respectively, of the domain boundary $\partial\Omega$ with $\partial\Omega=\partial_{\Psi}\Omega\cup\partial_{T}\Omega$ and $\partial_{\Psi}\Omega\cap\partial_{T}\Omega=\emptyset$, and N~N denotes the unit vector normal to $\partial\Omega$.
@@
$\int_{I}\left[\int_{\Omega}\left(\mathrm{Div}\,~P+\rho_{0}~B-\rho_{0}\ddot{\boldsymbol{\Psi}}\right)\cdot\boldsymbol{\xi}\,d\Omega+\int_{\partial_{T}\Omega}\left(~P~N\right)\cdot\boldsymbol{\xi}\,dS\right]\,dt=0,$ (3)
where $\boldsymbol{\xi}$ is a test function defined so that $\mathcal{V}\coloneqq\{\boldsymbol{\xi}\in W_{2}^{1}(\Omega\times I):\boldsymbol{\xi=~0}\text{ on }\partial_{\Psi}\Omega\times I\ \cup\Omega\times t_{0}\cup\Omega\times t_{f}\}$. The variational form in (3) is discretized in space using the classical Galerkin finite element method (FEM) [22]. The Neumann boundary condition from (2) is implemented by inserting PN=T~P~N=~T into the second integral in (3).
@@
$M\ddot{~u}+~K~u=~F$ (4)
where M~M is the (consistent) mass matrix, K~K is the stiffness matrix, F~F is the vector of applied external forces, $u=\boldsymbol{\Psi}(~X,t)-~X$ is the displacement, and $\ddot{~u}$ is the acceleration.
@@ 2.1 Model problem: wave propagation in a 1D linear elastic bar
$u(x,t_{0})=\frac{a}{2}\mathrm{exp}\left[-\frac{(x-b)^{2}}{2s^{2}}\right]$ (5)
for $a,b,s\in\mathbb{R}$.
To benchmark the performance of each transmission condition for NO-SAM, we developed a Python script that implements the numerical framework described above. All numerical examples in this work use the following problem setting, unless otherwise specified. A linear elastic constitutive model is employed with material constants E=1GPaE=1\,GPa and $\rho=1000\,kg\,m^{-3}$. The bar is of length L=1mL=1\,m and is discretized into uniform elements of length h=0.001mh=0.001\,m. The domain decomposition of the bar is defined as $\Omega_{1}=[0,0.6]$ and $\Omega_{2}=[0.6,1]$. The semi-discrete equation (4) is advanced in time using the Newmark-$\beta$ time integration scheme of constant average acceleration, for which $\beta=0.25$ and $\gamma=0.5$. The simulation is run from initial time $t_{0}=0\,s$ to final time $t_{f}=10^{-3}\,s$ using a time step of $\Delta t=2.5\times 10^{-7}\,s$. This time step was chosen based on the wave propagation speed $c=\sqrt{\frac{E}{\rho}}=1000\,m\,s^{-1}$, for which the stable time step was calculated to be $\Delta t_{stable}=\frac{h}{c}=10^{-6}\,s$ [34, 1]. Uniform, conformal meshes and identical time-step sizes are used for each subdomain, although SAM does not require this restriction. The initial displacement is defined using the symmetric Gaussian wave in (5) with $a=0.01$, $b=\frac{L}{2}=0.5$, and $s=0.02$.
@@ 3.1 NO-SAM formulation
As described in more detail in [34], the dynamic SAM algorithm operates by discretizing the temporal domain into time intervals $I_{n}=\left[t_{n},t_{n+1}\right]$ and solving a sequence of subdomain-local PDEs to convergence within each time interval. Specifically, for a given time interval $I_{n}$, NO-SAM solves the following sequence of subdomain-local problems:
$\left\{\begin{array}[]{rcll}~M_{1}\ddot{~u}_{1}^{(s+1)}+~K_{1}~u_{1}^{(s+1)}&=&~F_{1}&\text{in }\Omega_{1}\\[4.30554pt] ~u_{1}^{(s+1)}&=&~u_{D}&\text{on }\partial_{\Psi}\Omega_{1}\backslash\Gamma\\[2.15277pt] \alpha_{12}~T_{1}^{(s+1)}+\beta_{12}~u_{1}^{(s+1)}&=&~\lambda^{(s+1)}_{1}&\text{on }\Gamma\end{array}\right$. (6)
$\left\{\begin{array}[]{rcll}~M_{2}\ddot{~u}_{2}^{(s+1)}+~K_{2}~u_{2}^{(s+1)}&=&~F_{2}&\text{in }\Omega_{2}\\[4.30554pt] ~u_{2}^{(s+1)}&=&~u_{D}&\text{on }\partial_{\Psi}\Omega_{2}\backslash\Gamma\\[2.15277pt] \alpha_{21}~T_{2}^{(s+1)}+\beta_{21}~u_{2}^{(s+1)}&=&~\lambda^{(s+1)}_{2}&\text{on }\Gamma\end{array}\right$. (7)
$~\lambda^{(s+1)}_{1}$ $=\theta_{1}\left[\alpha_{12}~T_{2}^{(s)}+\beta_{12}~u_{2}^{(s)}\right]+(1-\theta_{1})\,~\lambda^{(s)}_{1}$ (8)
$~\lambda^{(s+1)}_{2}$ $=\theta_{2}\left[\alpha_{21}~T_{1}^{(s+1)}+\beta_{21}~u_{1}^{(s+1)}\right]+(1-\theta_{2})\,~\lambda^{(s)}_{2}$ (9)
where $s$ denotes the Schwarz iteration index and $\theta_{1},\theta_{2}\in(0,1]$ are relaxation parameters. For convenience, here we rewrite the Dirichlet boundary condition defined in (2) in terms of displacement, denoting the boundary condition on the subdomain interface as $u_{D}$. To allow for a general choice of transmission conditions, we introduce constants $\alpha_{ij},\beta_{ij}\in\mathbb{R}$ for $i,j=1,\dots,N_{k}$, where $N_{k}$ is the total number of subdomains ($N_{k}=2$ in the present study). The reader can observe that, by appropriate selection of these constants, Dirichlet, Neumann, and Robin transmission conditions can all be represented in (6)–(9).
@@
$\frac{\big{\lVert}\left[~u^{(s+1)}-~u^{(s)}\right]+\Delta t\,\left[~v^{(s+1)}-~v^{(s)}\right]+\tfrac{1}{2}\,\Delta t^{2}\,\left[~a^{(s+1)}-~a^{(s)}\right]\big{\rVert}_{2}}{\big{\lVert}~u^{(s)}+\Delta t\,~v^{(s)}+\tfrac{1}{2}\,\Delta t^{2}\,~a^{(s)}\big{\rVert}_{2}}<\delta$ (10)
@@
We consider formulations with these specific transmission conditions, as they have shown to lead to a convergent NO-SAM iteration in [47] and [29], respectively. Relaxation is typically introduced only to subdomains using a Dirichlet transmission condition on the boundary $\Gamma$, as occurs in the alternating Dirichlet-Neummann variant of NO-SAM, where $\alpha_{ij}=0$ and $\beta_{ij}\neq 0$. It has been demonstrated in the literature that, for certain PDEs, setting $\theta_{i}<1$ can reduce the number of Schwarz iterations required for convergence [47]. In this work, $\lambda^{i}_{0}$ is initialized to zero for the first iteration, i.e., $\lambda^{i}_{0}=~0$.
For the case of the Robin-Robin NO-SAM variant, the Robin transmission condition is applied as follows. Consider without loss of generality the subdomain problem in $\Omega_{1}$. In this case, the third line of (6) is rearranged to solve for $T_{1}^{(s+1)}$, and the resulting expression is then substituted into the external force term $F_{1}$ of the first line of (6), yielding
$M_{1}\ddot{~u}_{1}^{(s+1)}+\left[~K_{1}+\frac{\beta_{12}}{\alpha_{12}}~S_{1}\right]~u_{1}^{(s+1)}=~F_{1}+\frac{1}{\alpha_{12}}~R_{1}$ (11)
where the additional S~S and R~R terms can be viewed as Robin contributions to the stiffness matrix and the external force vector, respectively. The derivation and resulting expression for $\Omega_{2}$ is analogous, and is therefore omitted for brevity.

### v2

@@ Abstract
This work investigates transmission conditions for the domain decomposition-based coupling of subdomain-local models using the non-overlapping Schwarz alternating method (NO-SAM). Building on prior efforts involving overlapping SAM (O-SAM), we formulate and assess two NO-SAM variants, based on alternating Dirichlet–Neumann and Robin–Robin transmission conditions. For the subdomain-local models, we consider a mix of full order models (FOMs) and non-intrusive reduced order models (ROMs) constructed via an emerging model reduction technique known as operator inference (OpInf). Of particular novelty is the first application of NO-SAM to couple non-intrusive OpInf ROMs with each other, and with FOMs. Numerical studies on a one-dimensional linear elastic wave propagation benchmark problem demonstrate that transmission condition choice and parameter tuning significantly impact convergence rate, accuracy, and stability. Robin–Robin coupling often yields faster convergence than alternating Dirichlet–Neumann, though improper parameter selection can induce spurious oscillations at subdomain interfaces. For FOM–OpInf and OpInf–OpInf couplings, sufficient modal content in the ROM basis improves accuracy and mitigates instability, in some cases outperforming the coupled FOM–FOM reference solutions in both accuracy and efficiency. These findings highlight NO-SAM’s potential for enabling flexible, non-intrusive, and efficient multi-model coupling across independently meshed subdomains, while emphasizing the need for careful interface condition design in higher-dimensional and predictive settings.
@@ 1 Introduction
The present work advances a computational technique that promises to mitigate both hurdles described above: the Schwarz alternating method (SAM) [39]. The key idea behind SAM is to decompose the physical domain on which a given partial differential equation (PDE) is posed into smaller subdomains, and then to solve the governing PDE by iterating between smaller, subdomain-local problems, while exchanging boundary condition information to ensure compatibility across subdomain interfaces. Over the past decade, SAM has been shown to enable concurrent domain decomposition- (DD-)based coupling of subdomain-local full order models (FOMs) with each other [32, 33, 31], as well as with a variety of data-driven surrogate models (e.g., intrusive projection-based ROMs [46, 1], physics-informed neural networks (PINNs) [42], non-intrusive operator inference (OpInf) models [30]). In these papers, SAM exhibits a number of advantages over competing multi-scale coupling methods. Most notably: it is concurrent and minimally-intrusive, can couple non-conformal meshes with different element topologies, and allows each subdomain to be advanced with different time integrators and time steps for dynamic problems. Importantly, SAM achieves this without introducing non-physical artifacts into the solution.
@@
The work described here is an addition to a growing body of literature on DD-based OpInf-OpInf and ROM-FOM couplings. Our SAM approach is most similar to the recent work by Farcas et al. [13], which develops a method for performing DD-based coupling of subdomain-local OpInf ROMs by learning the appropriate reduced operators responsible for the coupling. The present work differs from that of Farcas et al. in two important ways: (i) in [13], each subdomain problem is solved once rather than by performing an iteration to convergence as done within our Schwarz framework, which may be insufficient for certain classes of nonlinear problems [34], and (ii) the approach in [13] is limited to overlapping domain decompositions (DDs). While other Schwarz-like methods for FOM-FOM and FOM-ROM couplings exist in the literature, the majority of these methods focus on the coupling of intrusive projection-based ROMs [6, 23, 24, 38] or neural network- (NN-)based ROMs [44, 45, 26, 27, 42]. A variety of other FOM-FOM and FOM-ROM coupling methods, e.g., Lagrange multipliers, optimization-based coupling, coupling via boundary maps, each with their own advantages and disadvantages, have also been proposed over the years. The interested reader is referred to [5, 7, 8, 22, 29, 36, 37, 11, 17, 3, 16, 41, 11, 18, 4, 43, 20, 12] and the references therein for more details. To the best of the authors’ knowledge, only two of these references, namely [16] and [13], consider the DD-based coupling of subdomain-local non-intrusive OpInf ROMs of the flavor considered herein [35].
@@ 2 Solid mechanics problem formulation
$\mathrm{Div}\,\mathord{\btensor P}+\rho_{0}\mathord{\btensor B}=\rho_{0}\ddot{\boldsymbol{\Psi}}\quad\text{in }\Omega\times I$ (1)
where $\Omega\in\mathbb{R}^{d}$ for $d\in\{1,2,3\}$ is an open bounded domain, $I\coloneqq\{t\in[t_{0},t_{f}]\}$ is a closed time interval with $0\leq t_{0}<t_{f}$, and $\mathord{\btensor x}=\boldsymbol{\Psi}(\mathord{\btensor X},t):\Omega\times I\rightarrow\mathbb{R}^{d}$ describes the mapping from the reference configuration to the current configuration with $\mathord{\btensor X}\in\Omega$. The overdot notation in $\ddot{\boldsymbol{\Psi}}$ denotes a time derivative such that $\dot{\boldsymbol{\Psi}}\coloneqq\frac{\partial\boldsymbol{\Psi}}{\partial t}$ and $\ddot{\boldsymbol{\Psi}}\coloneqq\frac{\partial^{2}\boldsymbol{\Psi}}{\partial t^{2}}$. The symbol $\mathord{\btensor P}$ denotes the first Piola-Kirchoff stress, $\mathord{\btensor B}$ is the specific body force, and $\rho_{0}$ is the mass density defined in the reference configuration. It is important to note that the constitutive model used to describe $\mathord{\btensor P}$ is general and can include nonlinear models.
@@
$\boldsymbol{\Psi}(\mathord{\btensor X},t_{0})$ $=\mathord{\btensor X}_{0}$ $\text{in }\Omega$ (2)
$\dot{\boldsymbol{\Psi}}(\mathord{\btensor X},t_{0})$ $=\mathord{\btensor v}_{0}$ $\text{in }\Omega$
$\boldsymbol{\Psi}(\mathord{\btensor X},t)$ $=\boldsymbol{\chi}$ $\text{on }\partial_{\Psi}\Omega\times I$
$\mathord{\btensor P}\mathord{\btensor N}$ $=\mathord{\btensor T}$ $\text{on }\partial_{T}\Omega\times I$
where $\partial_{\Psi}\Omega$ and $\partial_{T}\Omega$ denote the Dirichlet and Neumann portions, respectively, of the domain boundary $\partial\Omega$ with $\partial\Omega=\partial_{\Psi}\Omega\cup\partial_{T}\Omega$ and $\partial_{\Psi}\Omega\cap\partial_{T}\Omega=\emptyset$, and $\mathord{\btensor N}$ denotes the unit vector normal to $\partial\Omega$.
@@
[…]
@@
$\mathord{\btensor M}\ddot{\mathord{\btensor u}}+\mathord{\btensor K}\mathord{\btensor u}=\mathord{\btensor F}$ (4)
where $\mathord{\btensor M}\in\mathbb{R}^{N\times N}$ is the (consistent) mass matrix, $\mathord{\btensor K}\in\mathbb{R}^{N\times N}$ is the stiffness matrix, and $\mathord{\btensor F}\in\mathbb{R}^{N}$ is the vector of applied external forces. Here, $\mathord{\btensor u}\in\mathbb{R}^{N}$ is the vector of displacement degrees of freedom, defined as $\mathord{\btensor u}:=\boldsymbol{\Psi}(\mathord{\btensor X},t)-\mathord{\btensor X}$, and $\ddot{\mathord{\btensor u}}$ denotes the acceleration.
@@ 2.1 Model problem: wave propagation in a 1D linear elastic bar
$u(x,t_{0})=a\,\,\mathrm{exp}\left[-\frac{(x-b)^{2}}{2w^{2}}\right]$ (5)
for $a,b,w\in\mathbb{R}$, where these parameters represent the peak height, center, and standard deviation of the Gaussian wave, respectively.
To benchmark the performance of each transmission condition for NO-SAM, we developed a Python script that implements the numerical framework described above. All numerical examples in this work use the following problem setting, unless otherwise specified. A linear elastic constitutive model is employed with material constants E=1GPaE=1\,GPa and $\rho=1000\,kg\,m^{-3}$. The bar is of length L=1mL=1\,m and is discretized into uniform elements of length h=0.001mh=0.001\,m. The domain decomposition of the bar is defined as $\Omega_{1}=[0,0.6]$ and $\Omega_{2}=[0.6,1]$. The semi-discrete equation (4) is advanced in time using the Newmark-$\beta$ time integration scheme of constant average acceleration, for which $\beta=0.25$ and $\gamma=0.5$. The simulation is run from initial time $t_{0}=0\,s$ to final time $t_{f}=10^{-3}\,s$ using a time step of $\Delta t=2.5\times 10^{-7}\,s$. This time step was chosen based on the wave propagation speed $c=\sqrt{\frac{E}{\rho}}=1000\,m\,s^{-1}$, for which the stable time step was calculated to be $\Delta t_{stable}=\frac{h}{c}=10^{-6}\,s$ [33, 1]. Uniform, conformal meshes and identical time-step sizes are used for each subdomain, although SAM does not require this restriction. The initial displacement is defined using the symmetric Gaussian wave in (5) with a=0.005ma=0.005\,m, $b=\frac{L}{2}=0.5\,m$, and w=0.02mw=0.02\,m.
@@ 3.1 NO-SAM formulation
As described in more detail in [33], the dynamic SAM algorithm operates by discretizing the temporal domain into time intervals $I_{n}:=\left[t_{n},t_{n+1}\right]$ and solving a sequence of subdomain-local PDEs to convergence within each time interval. Specifically, for a given time interval $I_{n}$, NO-SAM solves the following sequence of subdomain-local problems:
$\left\{\begin{array}[]{rcll}\mathord{\btensor M}_{1}\ddot{\mathord{\btensor u}}_{1}^{(s+1)}+\mathord{\btensor K}_{1}\mathord{\btensor u}_{1}^{(s+1)}&=&\mathord{\btensor F}_{1}^{(s+1)}&\text{in }\Omega_{1}\\[4.30554pt] \mathord{\btensor u}_{1}^{(s+1)}&=&\mathord{\btensor u}_{D}&\text{on }\partial_{\Psi}\Omega_{1}\backslash\Gamma\\[2.15277pt] \alpha_{12}\mathord{\btensor T}_{1}^{(s+1)}+\beta_{12}\mathord{\btensor u}_{1}^{(s+1)}&=&{}^{(s+1)}_{1}&\text{on }\Gamma\end{array}\right$. (6)
$\left\{\begin{array}[]{rcll}\mathord{\btensor M}_{2}\ddot{\mathord{\btensor u}}_{2}^{(s+1)}+\mathord{\btensor K}_{2}\mathord{\btensor u}_{2}^{(s+1)}&=&\mathord{\btensor F}_{2}^{(s+1)}&\text{in }\Omega_{2}\\[4.30554pt] \mathord{\btensor u}_{2}^{(s+1)}&=&\mathord{\btensor u}_{D}&\text{on }\partial_{\Psi}\Omega_{2}\backslash\Gamma\\[2.15277pt] \alpha_{21}\mathord{\btensor T}_{2}^{(s+1)}+\beta_{21}\mathord{\btensor u}_{2}^{(s+1)}&=&{}^{(s+1)}_{2}&\text{on }\Gamma\end{array}\right$. (7)
${}^{(s+1)}_{1}$ $=\theta_{1}\left[\alpha_{12}\mathord{\btensor T}_{2}^{(s)}+\beta_{12}\mathord{\btensor u}_{2}^{(s)}\right]+(1-\theta_{1})\,{}^{(s)}_{1}$ (8)
${}^{(s+1)}_{2}$ $=\theta_{2}\left[\alpha_{21}\mathord{\btensor T}_{1}^{(s+1)}+\beta_{21}\mathord{\btensor u}_{1}^{(s+1)}\right]+(1-\theta_{2})\,{}^{(s)}_{2}$ (9)
where $s$ denotes the Schwarz iteration index and $\theta_{1},\theta_{2}\in(0,1]$ are relaxation parameters. For convenience, we rewrite the Dirichlet boundary condition defined in (2) in terms of displacement, denoting the boundary condition on the outer domain boundary as $\mathord{\btensor u}_{D}$. At each iteration, the displacements $\mathord{\btensor u}^{(s+1)}$ are solved for, and the solution on the interface $\Gamma$ is used to define the transmission conditions for the neighboring subdomain. To allow for a general choice of transmission conditions, we introduce constants $\alpha_{ij},\beta_{ij}\in\mathbb{R}$ for $i,j=1,\dots,N_{k}$, where $N_{k}$ is the total number of subdomains ($N_{k}=2$ in the present study). The reader can observe that, by appropriate selection of these constants, Dirichlet, Neumann, and Robin transmission conditions can all be represented in (6)–(9).
@@
$\frac{\big\lVert\left[\mathord{\btensor u}^{(s+1)}-\mathord{\btensor u}^{(s)}\right]+\Delta t\,\left[\mathord{\btensor v}^{(s+1)}-\mathord{\btensor v}^{(s)}\right]+\tfrac{1}{2}\,\Delta t^{2}\,\left[\mathord{\btensor a}^{(s+1)}-\mathord{\btensor a}^{(s)}\right]\big\rVert_{2}}{\big\lVert\mathord{\btensor u}^{(s)}+\Delta t\,\mathord{\btensor v}^{(s)}+\tfrac{1}{2}\,\Delta t^{2}\,\mathord{\btensor a}^{(s)}\big\rVert_{2}}<\delta$ (10)
@@
[…]
For the case of the Robin-Robin NO-SAM variant, the Robin transmission condition is applied as follows. Consider without loss of generality the subdomain problem in $\Omega_{1}$. In this case, the third line of (6) is rearranged to solve for $\mathord{\btensor T}_{1}^{(s+1)}$, and the resulting expression is then substituted into the external force term $\mathord{\btensor F}_{1}^{(s+1)}$ of the first line of (6), yielding
$\mathord{\btensor M}_{1}\ddot{\mathord{\btensor u}}_{1}^{(s+1)}+\left[\mathord{\btensor K}_{1}+\frac{\beta_{12}}{\alpha_{12}}\mathord{\btensor S}_{1}\right]\mathord{\btensor u}_{1}^{(s+1)}=\mathord{\btensor F}_{1}^{(s+1)}+\frac{1}{\alpha_{12}}\mathord{\btensor R}_{1}^{(s+1)}$ (11)
where the additional $\mathord{\btensor S}$ and $\mathord{\btensor R}$ terms can be viewed as Robin contributions to the stiffness matrix and external force vector, respectively. Their derivation is provided in the Appendix. The corresponding derivation for $\Omega_{2}$ is analogous and omitted for brevity.